PolHe3 Punch List For Restart

From HallCWiki
Jump to: navigation, search

Onsite Personnel Guidance

NOTE: You must follow the new MEDCON rules/etc noted on the sign-up page.
** Shift signup page
** Confirm travel plans in advance to help minimize 2 week quarantine on arrival! **
** Make sure your training and Site Visit status is still current to avoid delays. **

Training / Resumption of Physics Guidance

People

  • Stationed here w/ polhe3 as a priority
    • Brad, JP, Junhao, Mingyu, Bill, Arun
    • David Flay is new staff hire -- did thesis on 6 GeV d2n measurement! (JLab won't let him move to NN yet... grr)
    • Carlos Y (new Hall C postdoc) will be able to help by June
    • Mike Nycz (former Hall A grad student, now Temple post-doc) -- available to help
  • Offsite, but not too far away
    • [W&M] Todd, + 2 grad students (for shifts), post-doc (can RC)
    • [UVa] Xiaochao, Jixie, Gordon?, Chris J
  • Offsite, but plan to return when we know dates
    • Murchhana
    • Melanie

Tasks to complete before we get back on-site

  • These will likely require OSPs to be written up and approved
    • Remove/replace target enclosure panels
    • Laser/optics work on the target platform
    • Laser/optics work in the CH laser room (?)
    • Mounting a new cell (not an immediate task) [RadCon involvement too]
    • Install collimator for 1-pass running
    • EPR photo diode replacement?

Pre-beam Target Work

  • 'Spot check' B-field at target location for SHMS [Bill, Arun, Jixie]
    • 18 deg, -5.6 GeV/c as a cross check of the TOSCA modelling for production
    • 8.5 deg, -2.1286 GeV/c (electron QE) as a cross check of the TOSCA modelling for elastics
  • Test Kepco in current mode [Bill]
    • New power supply is now on-site at JLab
    • New filter (for 180 Hz) to install for current mode operation fix. Can always go to voltage mode if it doesn't work.
    • Establish/verify a known-good production mode
    • Further Details can be found here [1]
  • Update Target Cell documentation on wiki [Chris J, Arun, Mingyu]
    • Pol_He-3_Target_Information#Production_Target_Cell_Information
    • Many updates, not 100% sure if it is complete yet though (?)
      • Need to add supporting documentation/notes on where (and how) the measurements were taken
      • Need to add (and annotate) any duplicate/redundant measurements taken in different Labs (UVa, W&M, JLab)
  • pNMR issues [Mingyu]
    • Working, but some questions on errors and stability
    • Signal present in Transverse mode, but needs to be calibrated/cross checked when NMR works
    • pNMR Online analysis: Using Labview program to analysis pNMR FID signal (see [2])
      • Need to create a archived EPICS variable to record polarization measurement results from pNMR. [Brad, Chris, Mingyu]
  • Target work
    • Spin up curves for all 4 polarization directions, configurations

  • Laser Repair [Junhao, JP]
    • Shipped to Raytum; returned on ___________________ ?
  • Purchase another reference cell [Todd, JP]
    • New reference cell 'Fauci' arrived and was part of the alignment tests in the Hall July 10. Unfortunately the copper joint broke (known risk)
    • Will follow up and get it repaired.
    • Repair status, PO is out
  • Moller cryo control issues [Gaskell]
    • JT valve controller isn't working right (27 Jul 2020)
  • Update Target Operator documentation [Arun, et al.]
    • Double check that all EPICS variables are working!
    • Ensure all NMR, EPR calibration constants are documented and updated in software
  • Locate W-shielding block [JP, Dave H.]
    • Found [3]
    • Back in the roped area to beam-right of the HMS
  • Follow up Walter on the compressor water issue
  • Modify optics foil configuration (improve HMS optics) [Arun, RadCon]
    • See HCLOG 3751348 for diagram.
    • Now: -20, 0, 13.34, 20.  : Current configuration has foils 1, 4*, 6*, 7, where (*) indicates a hole target.
    • Want: -13.34, -6.67, 0, 6.67, 13.34, 20. : Foils 2, 3, 4*, 5, 6*, 7 (Remove 1; Replace 2, 3, 5)
    • Updated: https://logbooks.jlab.org/entry/3816056
  • We have agreed to rotate the HMS to 20 degrees (kin C) and start running with the HMS at that angle. (The SHMS will remain at its current 11 degree 'Kin X'.) This will allow the SOLID Prototype folks to get their small-angle test completed with minimal impact on our production running. (They will setup at small angle and do all their checkout tests well before beam.)
    • I (Brad) think this sounds OK, but everyone should think through what we need to do and let us know if there is anything that would require moving the HMS in to small angles, even temporarily!, during our pre-beam setup!
  • EPR problems 
    [4] saw some progress on Tuesday [5], but issues remain.
    • Tentative plan now is to continue the d2n program by replacing the photo-diode that is placed directly on the pumping chamber (we have many spares).
      • It would be good to write up a procedure to streamline this replacement procedure for us and RadCon
    • Maybe design a new mount for the photodiode replacing the old mount designed for the fiber bundle
  • NMR issues [Junhao]
    • Transverse NMR is now working
    • NMR Calibration Constants : Verify that calibration constants and input values associated with NMR and other polarimetry are consistent and will be correct when the software is restarted
  • Restore items noted in HCLOG 3809218 when site access is restored. [Brad]
  • Spectrum analyzer for transverse configuration ASAP [Junhao]
    • Needs fiber pulled from target to behind
  • Laser Control Software Update [Carlin, Brad]
    • Three new lasers require modifications/updates to our existing laser control software.
      • These changes must be integrated with the existing EPICS logging that has been added to the our laser control software.
    • Chris Carlin has agreed to take care of this. He has the software to do the updates in parallel, but will need some testing time on polhe1 later this week. (29 Jun 2020)
  • AnywhereUSB bridge issues [Brad]
    • Contact with Target readbacks through AnywhereUSB USB bridge still drop out on occasion
    • Description of situation during A1n run:
      • At first, we fixed a network problem raised by Ethan. And this decreased the frequency of the drop out
      • Later, there are two scenarios for the drop out. For the first scenario which happened more frequently, after the device dropped out, we just need to close the individual labview programs with the devices dropped out and wait for the device to come back then reopen the program. If the device do not show up in the first scenario, then we need to close the Labview and restart the computer. (Seems like more of a software issue, not hardware?)

Pending Issues

  • 1-pass running considerations [Brad, Lassiter, Mark]
    • HMS setting: 11.7 deg, -2.1483 GeV/c; SHMS setting: 8.5 deg, -2.12860 GeV/c
    • Install collimator
    • Turn off hodoscope bars as before: HMS [11]; SHMS [12]
    • Last time there was an issue with the raster; make sure we know what needs to be done [13] [Mark]
    • Lots of good notes here: 1-pass_calibration_for_A1n_(Jan_2020)
  • Extended-target optics analysis [Mark, Holly]
    • Incorporate into online (and offline) replay code (**)
    • SHMS optics is in good shape; HMS pretty good, but could use a more 16.4deg data with additional foils (see note in first section).
    • HMS could use more data at 16.5 degree setting

  • Online analysis software ready to go? [Sylvester?, Brad]
    • Most should be fine, but are we missing any 'physics' plots?
    • Merge and test code updates to hcana / hallc_replay:
      • Better error handling and typo in our SHMS shower PARAM file [14]
  • 2mm reference cell offset [Todd, JP, Brad]
    • Is there a way to shift the reference cell position before beam starts?
    • Arun's HCLOG 3807201 about the unfortunate reference cell alignment vs. the rest of the target ladder.
    • NOTE: "Shift Leader must request a beam that is shifted horizontally to the beam left side by 2 mm to prevent scraping."
  • Ensure spectrometer rotation is ready to go
    • HMS: should be parked at 20deg and rotation disabled (no rotation allowed until SOLID test stand is removed)
    • SHMS: confirm rotation between 18--11deg (5-pass production range); rotate to 11.0 deg for start up
      • Remind that we will be moving both spectrometers to forward angles for 1-pass running
  • Verify that all other EPICS channels are present and are still being archived [Brad]
    • Seems to be OK -- leaving this here to recheck just before we start
  • Get 3H07B Harp repaired [15] [16]; and a puzzle [17]
    • Repair work is now scheduled (nominal start June 22) [18]
      • Harp reinstalled week of July 13; final alignment to be done on Monday July 20
  • There is a remaining issue with a replacement scaler [19] [Steve, Brad]
    • Steve will follow up on this Monday/Tuesday
    • Verify that inputs to 'Scale32'/G0 scalers (ie ROC5) are > 10ns in width. [Brad, Steve]
  • Verify SHMS, HMS magnets are ready to go [Brad, Lassiter]
  • Look at trigger related signals on the scope under high load (reftime extra pules, FADC reftime-shaper pileup?, NGC noise/rate questions) [Brad]
    • Should be corrected [20]
  • Investigate FADC 'blocking' artifact (pulse pile-up, undershoot, weirdness?) for reftime shapers
    • Should be corrected [21]
  • Check cabling on SHMS Shower Layer1, PMT4 timing (25ns difference vs other PMTs) [Brad]
    • Didn't see anything obvious, will check again
    • No issues seen [22]
  • Check cabling on NGC PMT3 (12 ns later than other PMTs in TDC spectrum) [Brad]
    • Added shim cable; resolved [23]
  • HMS Q3 LCW leak --supply and return valves were closed [24]
  • Fix the EPICS variable calculation for target field angle [Bill]
    • This is the PV which is calculated from the coils currents that is in the startOfRun entries.
    • It needs a new formula since we changes the coil orientation
  • Target Ladder software must throw an error on a 'voltage' readback mismatch! [25]
    • Ethan reports this is complete and an updated screen is installed
  • iochc10 issues [Gaskell]
    • iochc10 should be removed from OPS monitoring systems -- Hall C monitors it now (5/27 DG)
  • Compton related items in Hall C beam steer-up procedure still cause confusion between MCC and Shift Crew [Gaskell]
    • Procedures need to be updated at MCC (Should be completed 5/27 DG)

Pre-beam punchlist

Copied from pre-beam punchlist done just before MC6 shutdown. Some items should be completed but require verification/confirmation.

Techs w/ RadCon [Todd (or designate) for cooling jets]

  • Turn on turbos for downstream beamline
  • Turn on turbos for SHMS and HMS spectrometer vacuums
  • Remove the guards on the SHMS and HMS entrance windows (verify)
  • Remove the guards on the both upstream and downstream beamline windows (verify)
  • Ensure the beamline window cooling jets are in position and on (verify)
  • Ensure the target window cooling jets are in position and on (verify)
  •  ??

Miscellaneous

  • [Gaskell] Moller work and cool-down [26]
  • [RadCon, RC] RadCon checklist
  • [Gaskell] Beamline checklist
  • [Brad] Detector/DAQ checklist
  • [Walter] Hall checklist
  • [OPS, RC] Schedule hot checkout of Hall C arc magnets/power-supplies
  • [OPS, RC] Verify all relevant FSDs that may have been masked for the down are known about and ready to unmasked when appropriate
  • [Lassiter] Enable all spectrometer magnet supplies, restore production settings for Cryo
  • [Brad] Double check that rotation is enabled in the GUIs for both spectrometers
  • [Brad] Ramp all SHMS and HMS magnets to something modest (quick check), or to first kinematic. For example
    • SHMS: 11.0deg, 7.5 GeV/c
    • HMS: 13.5deg, 4.2 GeV/c
    • NOTE: Coordinate with the target group on the appropriate HB setting so we don't interfere with any target pumping.
  • [Techs, Brad?] Do a final check to ensure that there are no potential issues with remote spectrometer rotation
  • Restart main alarm handler (resets masked channels)
  • Restart target alarm handler (if needed)
  •  ??